# Eccentric Signatures of Stellar-Mass Binary Black Holes with Circumbinary Disks in LISA

### v1

@@ Abstract
Binary black holes may have circumbinary disks if formed through common-envelope evolution or within gaseous environments. Disks can drive binaries into wider and more eccentric orbits, while gravitational waves harden and circularise them. We combine cutting-edge evolution prescriptions for disk-driven binaries with well-known equations for gravitational-wave-driven evolution, and study the evolution of stellar-mass binary black holes. We find that binaries are driven by their disk to an equilibrium eccentricity, $0.2\lesssim e_{\mathrm{eq}}\lesssim 0.5$, that dominates their evolution. Once they transition to the GW-dominated regime their eccentricity decreases rapidly; we find that binaries with long-lived disks will likely be observed in LISA with detectable eccentricities $\sim 10^{-2}$ at $0.01$ Hz, with the precise value closely correlating with the binary’s initial mass ratio. This may lead binary black holes with CBDs observed in LISA to be confused with dynamically-formed binary black holes.
@@ 1 Introduction
Binary black holes (BBHs) evolving within circumbinary disks (CBDs) can be driven to long-lasting equilibrium eccentricities (D’Orazio & Duffell, 2021; Zrake et al., 2021; Siwek et al., 2023b). Merging stellar-mass BBHs form via several pathways that may involve CBDs: for example, an isolated binary may go through common envelope (CE) evolution (Paczynski, 1976; Livio & Soker, 1988; Bethe & Brown, 1998; Dominik et al., 2012; Ivanova et al., 2013; Kruckow et al., 2016; Stevenson et al., 2017), while a dynamically-assembled binary may evolving while embedded in an active galactic nuclei (AGN) disk (Bartos et al., 2017; McKernan et al., 2018; Yang et al., 2019; McKernan et al., 2020; Gröbner et al., 2020; Ford & McKernan, 2022; Samsing et al., 2022; Calcino et al., 2023), and binaries evolving in field triples may also gain CBDs (Silsbee & Tremaine, 2017; Antonini et al., 2017a; Liu & Lai, 2017; Liu et al., 2019; Vigna-Gómez et al., 2021; Dorozsmai et al., 2024). Dynamical assembly in dense star clusters is not thought to produce BBHs with CBDs, since frequent BH partner-swappings and interactions likely disrupt disk formation (Wen, 2003; Askar et al., 2017; Rodriguez et al., 2018a, b; Samsing, 2018; Antonini & Rasio, 2016; Hoang et al., 2018; Fragione & Bromberg, 2019; Chattopadhyay et al., 2023). Identifying the influence of a CBD on the orbit of a BBH can therefore distinguish between different formation environments.
For BBHs detected with the current generation of ground-based detectors, measurable eccentricity is considered a robust indicator of dynamical formation (Lower et al., 2018; Zevin et al., 2021). Isolated BBHs are expected to have negligible eccentricities in ground-based detectors, but close-to-merger eccentricities in dynamically-formed or tertiary-driven BBH can be higher, with different pathways yielding distinct eccentricity distributions (e.g., Samsing, 2018; Antonini et al., 2017b; Rodriguez et al., 2018a; Tagawa et al., 2021; Arca Sedda et al., 2021; Samsing et al., 2022). The limited eccentricity sensitivity of Advanced LIGO-Virgo-KAGRA (LVK) (Abbott et al., 2018) detectors permits a narrow window onto these distributions: we are sensitive to only $e_{10}\gtrsim 0.05$ at GW frequencies of $10$ Hz (Lower et al., 2018), while simulations of dynamical environments predict distinct distributions with structures down to $e_{10}\sim 10^{-8}$.
@@
Since GW frequency increases as the binary separation shrinks, earlier stages of a BBH’s life will be visible to LISA than to ground-based detectors. Since GWs circularise orbits, younger systems are also likely to have higher eccentricities (Peters, 1964). With LISA sensitivity, eccentricities $\gtrsim 10^{-3}$ can be measured for BBHs inspiralling through the LISA band (Klein et al., 2022; Garg et al., 2023; Wang et al., 2024). Binaries evolving either in isolation or after being ejected from their host cluster, which have $e_{10}\sim 10^{-8}$, are expected to have $10^{-3}\lesssim e_{\mathrm{0.01}}\lesssim 0.9$ at $0.01$ Hz, and in-cluster mergers will have $e_{\mathrm{0.01}}\gtrsim 10^{-2}$ (DePorzio et al., 2024; Wang et al., 2024). Eccentricity is therefore likely to be a common feature of stellar-mass BBHs in LISA.
Interactions of a BBH with a CBD can cause distinct observables, which may facilitate the identification of CBD-driven BBH mergers. Such observables include electromagnetic counterparts (e.g., Mösta et al., 2010; Martin et al., 2018), and mass ratios closer to unity due to preferential accretion onto the less massive component (e.g., Farris et al., 2014; Gerosa et al., 2015; Duffell et al., 2020; Siwek et al., 2023a). As we show in this work, another feature that could distinguish BBHs with CBDs is distinct and measurable eccentricity values at $0.01$ Hz, $0.2\lesssim e_{0.01}\lesssim 0.5$. We find that the value of $e_{0.01}$ correlates with the binary’s initial mass ratio $q_{i}$, meaning that it may also correlate with the final mass ratio $q_{f}$ if the BBH torques the CBD sufficiently to prevent accretion (Martin et al., 2018). Alternatively, a population of $q_{f}\approx 1$ BBHs may have their initial distribution of $q_{i}$ mapped by their detected $e_{0.01}$.
@@
Ishibashi & Gröbner (2020) developed an analytic model to combine the driving factors of gas torques and GW emission on BBH embedded in AGN disks, with disk properties dependent on the mass and proximity of the central supermassive BH. They estimate the eccentricities of equal-mass BBH with $M=50$ M in LISA to be in the range 0.010.010.010.010.10.10.10.1. Ishibashi & Gröbner (2024) updated the simplified analytic model to include the effects of accretion, studying three different total masses and five different mass ratios. Zrake et al. (2021) combine hydrodynamical simulations, described in Tiede et al. (2020), with equations for GW-driven evolution from Peters (1964), to evolve equal-mass BBHs from stellar to supermassive, finding all equal-mass BBHs have $e_{\mathrm{eq}}\sim 0.45$.
@@ 2 Method
The gas-driven evolution of BBHs is encapsulated by incremental changes in semi-major axis, $\dot{a}_{\mathrm{gas}}$, and eccentricity, $\dot{e}_{\mathrm{gas}}$. Rates of change are derived in Siwek et al. (2023b) for systems with different initial mass ratios, $q_{i}$, and initial eccentricities, $e_{i}$. The evolution rates are scaled by the initial separation, $a_{i}$, and initial total mass of the binary, $M_{i}$. Binaries are evolved using a grid of hydrodynamical simulations, created using the Navier Stokes version of moving-mesh code Arepo (Springel, 2010; Pakmor et al., 2016). The disk viscosity, $v=\alpha(h/r)^{2}r^{2}\sqrt{GM/r^{3}}$, is parameterized by a constant $\alpha=0.1$ (Shakura & Sunyaev, 1973), and the disk aspect ratio is fixed at $h/r=0.1$. The results of these simulations are in good agreement with the results of simulations using other codes that employ similar methods, e.g., Muñoz et al. (2019); D’Orazio & Duffell (2021); Zrake et al. (2021).
The evolution rates calculated in Siwek et al. (2023b) are valid for generic binary-plus-disk systems. Their scaling to physical quantities depends on the ratio of accretion efficiency in Eddington units, $f_{e}$, to the radiative efficiency parameter $\epsilon$. The latter may take a value between 00 and $1$, and represents the fraction of the rest mass energy of the object lost as radiation during the accretion process. The gravitational potential energy of a particle in the disk can be lost due to friction-induced gas heating. For single BHs, $\epsilon$ ranges between about $0.06$ for Schwarzchild BHs and $0.32$ for maximally-spinning BHs (Laor & Netzer, 1989). Interactions of BHs with each other and with their circumbinary disk, as well as poorly-constrained mechanisms of radiative feedback, make $\epsilon$ uncertain for BBH evolution. We use a conservative estimate of $\epsilon=0.1$. We note that $f_{e}/\epsilon=0.1$ (when $f_{e}=0.01,\epsilon=0.1$) could equally be achieved with, for example, $f_{e}=0.1,\epsilon=1$.
@@
Figure 1: The eccentricity evolution of BBHs with $e_{i}=0.1,0.8$ (left and right columns, respectively, of each set of plots) and $q_{i}=0.1,0.5,1.0$ (top to bottom rows) with initial $m_{1,i}=30$ M. The left set of plots shows the eccentricity evolution with peak GW frequency, while the right set of plots shows eccentricity evolution with time. The assumed $f_{e}$ is indicated by the opacity of the curve. Because the CBD forces binaries with high $e_{i}$ to smaller eccentricities much more rapidly than it shrinks their separation, these binaries evolve from higher to lower peak GW frequency before reaching their steady-state. The steady-state eccentricity decays above $10^{-4}$ Hz.
@@
Figure 2: Eccentricities at $0.01$ Hz for all binaries with $e_{i}>0.05$. Square and circular scatter points correspond to $f_{e}=0.01$ and $0.1$ respectively. Black-edged markers correspond to our main results with unlimited accretion and red markers correspond to the negligible accretion case. The initial eccentricity corresponds to the alpha value of the scatter marker (higher $e_{i}$, higher opacity). For the negligible accretion case, $e_{i}$ minimally affects $e_{0.01}$, while for unlimited accretion lower $e_{i}$ can lead to slightly lower $e_{0.01}$. All values are detectable with LISA (Klein et al., 2022; Wang et al., 2024).
@@ 2.1 Accretion of gas
While $f_{e}$ represents the fraction of the Eddington accretion rate that a single BH of mass $M=m_{1}+m_{2}$ would accrete at, the individual BHs likely accrete asymmetrically at different rates (Siwek et al., 2023a). This parameter can be considered a proxy for the surface density of the disk, providing a measure of how much gas is available to torque the binary. The actual efficiency of mass accretion onto stellar-mass BBH from disks formed in post-CE systems is influenced by multiple uncertain factors, e.g., magnetic fields in gas disks of this type, which affect the rate of viscous transport. While we allow unlimited accretion in our main results, we compare these to two other cases. Firstly, we consider negligible accretion, since “typical” binary-disk parameters for post-CE stellar-mass BBH are found to converge towards the zero-accretion case due to the torques that the binary exerts on the disk Martin et al. (2018). Secondly, we compare to the limited-accretion cases, representing a disk with finite mass.
@@ 3 Results
BBHs with CBDs are candidates for multi-band observations and archival searches in LISA data. Fig. 2 shows the expected eccentricity at $0.01$ Hz, $e_{0.01}$. We find $e_{0.01}\sim 10^{-2}$, within the range of detectability for such searches, and of the same magnitude as expected eccentricities from dynamically-formed mergers (Wang et al., 2024). The $e_{0.01}$ values remain set primarily by $q_{i}$, although for the unlimited-accretion case we expect $q_{f}\approx 1$. For the negligible-accretion case, in which $q_{f}=q_{i}$, this correlation a potential smoking-gun of BBHs with CBDs.
@@
By the time the $f_{\mathrm{GW}}$ reaches $1$ Hz, the optimistic minimum frequency of the future ground-based detector Einstein Telescope (ET; Maggiore et al., 2020), BBHs with CBDs have eccentricities $e_{1}\approx 10^{-4}$ if $f_{e}=0.1$, equivalent to $e_{10}\approx 10^{-5}$, below the expected sensitivity for ET (Lower et al., 2018; Saini, 2024).
@@ 4 Discussion
The LISA-band eccentricity evolution of a stellar-mass BBH is markedly different when it has a CBD, as the CBD drives the binary to a characteristic equilibrium eccentricity determined by its initial mass ratio. We find that both $e_{\rm eq}$ and $e_{0.01}$ directly correspond to $q_{i}$, insensitively to assumptions about accretion efficiency. As long as the disk endures, the eccentricity at $f_{\mathrm{GW}}=0.01$ Hz is $e_{0.01}\sim 10^{-2}$, detectable for LISA and in the range usually expected for diskless dynamically-formed in-cluster mergers (Klein et al., 2022; Garg et al., 2023; Wang et al., 2024; DePorzio et al., 2024). Since $e_{0.01}$ is a strong function of $q_{i}$, a population of BBHs observed with an eccentricity-mass ratio correlation would indicate contributions from non-accreting BBHs with CBDs, while a population of equal-mass BBHs with eccentricities $\sim 2\times(10^{-3}$$10^{-2})$ would more likely come from accreting BBHs with CBDs.
@@
The eccentricity evolution of stellar-mass BBHs with CBDs is similar to that of the supermassive BBHs evolved in Siwek et al. (2024), but proceeds at higher frequencies and slightly higher eccentricities due to the lower mass. Our results using limited accretion agree with those of Valli et al. (2024): disks with $M_{d}=0.1M_{i}$ do not drastically influence the separation or mass ratio of the BBH, but do drive the binary eccentricity to its equilibrium value. The $e_{\mathrm{eq}}$ values we find for stellar-mass BBHs are in good agreement with the equilibrium eccentricity values found in Valli et al. (2024), with the maximum equilibrium eccentricity of $e_{\mathrm{eq}}\approx 0.5$ occurring for $q_{i}=0.7$. However, in contrast to previous work including Siwek et al. (2023b, a); Valli et al. (2024), none of our binaries undergoes orbital expansion, since $\dot{a}_{\mathrm{GW}}\gg\dot{a}_{\mathrm{gas}}$ when $a_{i}=0.2$ au.

### v3

@@ Abstract
Stellar-mass binary black holes may have circumbinary disks if formed through common-envelope evolution or within gaseous environments. Disks can drive binaries into wider and more eccentric orbits, while gravitational waves harden and circularise them. We combine cutting-edge evolution prescriptions for disk-driven binaries with well-known equations for gravitational-wave-driven evolution, and study the evolution of stellar-mass binary black holes. We find that binaries are driven by their disk to an equilibrium eccentricity, $0.2\lesssim e_{\mathrm{eq}}\lesssim 0.5$, that dominates their evolution. Once they transition to the GW-dominated regime their eccentricity decreases rapidly; we find that stellar-mass binary black holes with long-lived disks will likely be observed in LISA with detectable eccentricities $\sim 10^{-2}$ at $0.01$ Hz, with the precise value closely correlating with the binary’s initial mass ratio. This may lead stellar-mass binary black holes with CBDs observed in LISA to be confused with dynamically-formed binary black holes.
@@ 1 Introduction
Binary black holes (BBHs) evolving within circumbinary disks (CBDs) can be driven to long-lasting equilibrium eccentricities (D’Orazio & Duffell, 2021; Zrake et al., 2021; Siwek et al., 2023b). Merging stellar-mass BBHs form via several pathways that may involve CBDs: for example, an isolated binary may go through common envelope (CE) evolution (Paczynski, 1976; Livio & Soker, 1988; Bethe & Brown, 1998; Dominik et al., 2012; Ivanova et al., 2013; Kruckow et al., 2016; Stevenson et al., 2017), while a dynamically-assembled binary may evolve while embedded in an active galactic nuclei (AGN) disk (Bartos et al., 2017; McKernan et al., 2018; Yang et al., 2019; McKernan et al., 2020; Gröbner et al., 2020; Ford & McKernan, 2022; Samsing et al., 2022; Calcino et al., 2023), and binaries evolving in field triples may also gain CBDs (Silsbee & Tremaine, 2017; Antonini et al., 2017a; Liu & Lai, 2017; Liu et al., 2019; Vigna-Gómez et al., 2021; Dorozsmai et al., 2024). Dynamical assembly in dense star clusters is not thought to produce BBHs with CBDs, since frequent BH partner-swappings and interactions likely disrupt disk formation (Wen, 2003; Askar et al., 2017; Rodriguez et al., 2018a, b; Samsing, 2018; Antonini & Rasio, 2016; Hoang et al., 2018; Fragione & Bromberg, 2019; Chattopadhyay et al., 2023). Identifying the influence of a CBD on the orbit of a BBH can therefore distinguish between different formation environments.
For BBHs detected with the current generation of ground-based detectors, measurable eccentricity is considered a robust indicator of dynamical formation (Lower et al., 2018; Zevin et al., 2021). Isolated BBHs are expected to have negligible eccentricities in ground-based detectors, but close-to-merger eccentricities in dynamically-formed or tertiary-driven BBH can be higher, with different pathways yielding distinct eccentricity distributions (e.g., Samsing, 2018; Antonini et al., 2017b; Rodriguez et al., 2018a; Tagawa et al., 2021; Arca Sedda et al., 2021; Samsing et al., 2022). For example, BBHs forming in globular clusters (GCs) have $10^{-8}\lesssim e_{10}\lesssim 10^{-3}$ if ejected from the cluster, $10^{-7}\lesssim e_{10}\lesssim 10^{-2}$ if they merge between dynamical interactions, and $10^{-3}\lesssim e_{10}\lesssim 1$ if they merge through GW capture (Zevin et al., 2021). The limited eccentricity sensitivity of Advanced LIGO-Virgo-KAGRA (LVK) (Abbott et al., 2018) detectors permits a narrow window onto eccentricity distributions: we are sensitive to only $e_{10}\gtrsim 0.05$ at GW frequencies of $10$ Hz (Lower et al., 2018). Just a small fraction of the multi-peaked distribution predicted for BBHs in GCs, which differs from those arising in other environments like AGN disks and field triples, can be used to distinguish a population of GC mergers from other channels in current detectors (Romero-Shaw et al., 2022).
@@
Since GW frequency increases as the binary separation shrinks, LISA will observe earlier stages of a stellar-mass BBH’s life compared to ground-based detectors. Since GWs circularise orbits, younger systems are also likely to have higher eccentricities (Peters, 1964). With LISA sensitivity, eccentricities $\gtrsim 10^{-3}$ can be measured for stellar-mass BBHs inspiralling through the LISA band at a GW frequency of $0.01$ Hz (Klein et al., 2022; Garg et al., 2023; Wang et al., 2024). Binaries evolving in isolation are expected to have $10^{-5}\lesssim e_{\mathrm{0.01}}\lesssim 10^{-3}$ at $0.01$ Hz, those merging after ejection from their host cluster have $10^{-4}\lesssim e_{\mathrm{0.01}}\lesssim 10^{-2}$, and in-cluster mergers will have $e_{\mathrm{0.01}}\gtrsim 10^{-2}$ (Wang et al., 2024). Eccentricity is therefore likely to be a common feature of stellar-mass BBHs in LISA.
Interactions of a BBH with a CBD can cause distinct observables, which may facilitate the identification of CBD-driven BBH mergers. Such observables include electromagnetic counterparts (e.g., Mösta et al., 2010; Martin et al., 2018), and mass ratios closer to unity due to preferential accretion onto the less massive component (e.g., Farris et al., 2014; Gerosa et al., 2015; Duffell et al., 2020; Siwek et al., 2023a). As we show in this work, another feature that could distinguish stellar-mass BBHs with CBDs is distinct and measurable eccentricity values at $0.01$ Hz, $e_{0.01}\sim 10^{-2}$. We find that the value of $e_{0.01}$ correlates with the binary’s initial mass ratio $q_{i}$, meaning that it may also correlate with the final mass ratio $q_{f}$ if the BBH torques the CBD sufficiently to prevent accretion (Martin et al., 2018). Alternatively, a population of $q_{f}\approx 1$ BBHs may have their initial distribution of $q_{i}$ mapped by their detected $e_{0.01}$.
@@
Ishibashi & Gröbner (2020) developed an analytic model to combine the driving factors of gas torques and GW emission on BBH embedded in AGN disks, with disk properties dependent on the mass and proximity of the central supermassive BH. They estimate the eccentricities of equal-mass BBH with $M=50$ M in LISA to be in the range 0.010.010.010.010.10.10.10.1. Ishibashi & Gröbner (2024) updated the simplified analytic model to include the effects of accretion, studying three different total masses and five different mass ratios. Zrake et al. (2021) combine hydrodynamical simulations with equations for GW-driven evolution from Peters (1964) to evolve equal-mass BBHs of masses ranging from stellar to supermassive, finding all equal-mass BBHs have $e_{\mathrm{eq}}\sim 0.45$as a result of gas driving, and that an $M=147$ M binary at redshift $z=1$ has $e\sim 5\times 10^{-3}$ at $\sim 0.01$ Hz. For other classes of LISA sources, such as the much heavier and louder supermassive BBH mergers, eccentricities can be measured at lower frequencies; for example, Armitage & Natarajan (2005); Roedig et al. (2011); Siwek et al. (2024) all predict that $10^{6}$ M BBHs retain $e\gtrsim 10^{-3}$ in LISA.
@@ 2 Method
The gas-driven evolution of BBHs is encapsulated by incremental changes in semi-major axis, $\dot{a}_{\mathrm{gas}}$, and eccentricity, $\dot{e}_{\mathrm{gas}}$. Rates of change are derived in Siwek et al. (2023b) for systems with different initial mass ratios, $q_{i}$, and initial eccentricities, $e_{i}$. The evolution rates are scaled by the initial separation, $a_{i}$, and initial total mass of the binary, $M_{i}$. We use the orbital evolution prescriptions published in Siwek et al. (2023b), computed using a grid of hydrodynamical simulations created using the Navier Stokes version of moving-mesh code Arepo (Springel, 2010; Pakmor et al., 2016). The disk viscosity, $v=\alpha(h/r)^{2}r^{2}\sqrt{GM/r^{3}}$, is parameterized by a constant $\alpha=0.1$ (Shakura & Sunyaev, 1973), and the disk aspect ratio is fixed at $h/r=0.1$. The results of these simulations are in good agreement with the results of simulations using other codes that employ similar methods, e.g., Muñoz et al. (2019); D’Orazio & Duffell (2021); Zrake et al. (2021).
The evolution rates calculated in Siwek et al. (2023b) are valid for generic binary-plus-disk systems. Their scaling to physical quantities, e.g., the absolute accretion rate and electromagnetic luminosity, depends on the ratio of accretion efficiency in Eddington units, $f_{e}$, to the radiative efficiency parameter $\epsilon$. The latter may take a value between 00 and $1$, and represents the fraction of the rest mass energy of the object lost as radiation during the accretion process. The gravitational potential energy of a particle in the disk can be lost due to friction-induced gas heating. For single BHs, $\epsilon$ ranges between about $0.06$ for Schwarzchild BHs and $0.32$ for maximally-spinning BHs (Laor & Netzer, 1989). Interactions of BHs with each other and with their circumbinary disk, as well as poorly-constrained mechanisms of radiative feedback, make $\epsilon$ uncertain for BBH evolution. We use a conservative estimate of $\epsilon=0.1$. We note that $f_{e}/\epsilon=0.1$ (when $f_{e}=0.01,\epsilon=0.1$) could equally be achieved with, for example, $f_{e}=0.1,\epsilon=1$.
@@
Figure 1: The eccentricity evolution of BBHs with $e_{i}=0.1,0.8$ (left and right columns, respectively, of each set of plots) and $q_{i}=0.1,0.5,1.0$ (top to bottom rows) with initial $m_{1,i}=30$ M, assuming unlimited accretion (see Sec. 2.1). The left set of plots shows the eccentricity evolution with peak GW frequency, while the right set of plots shows eccentricity evolution with time. The assumed $f_{e}$ is indicated by the opacity of the curve. Because the CBD forces binaries with high $e_{i}$ to smaller eccentricities much more rapidly than it shrinks their separation, these binaries evolve from higher to lower peak GW frequency before reaching their steady-state. The steady-state eccentricity decays above $10^{-4}$ Hz.
@@
Figure 2: Eccentricities at $0.01$ Hz for all binaries with $e_{i}>0.05$. Square and circular scatter points correspond to $f_{e}=0.01$ and $0.1$ respectively. Colours correspond to the cases of unlimited acccretion and no accretion ("none", red markers), both with long-lived disks; binaries without disks have $e_{0.01}<10^{-3}$. The initial eccentricity corresponds to the alpha value of the scatter marker (higher $e_{i}$, higher opacity). For the negligible accretion case, $e_{i}$ minimally affects $e_{0.01}$, while for unlimited accretion lower $e_{i}$ can lead to slightly lower $e_{0.01}$. All values are detectable with LISA (Klein et al., 2022; Wang et al., 2024).
@@ 2.1 Accretion of gas
While $f_{e}$ represents the fraction of the Eddington accretion rate that a single BH of mass $M=m_{1}+m_{2}$ would accrete at, the individual BHs likely accrete asymmetrically at different rates (Siwek et al., 2023a). This parameter can be considered a proxy for the surface density of the disk, providing a measure of how much gas is available to torque the binary: the denser the disk, the stronger the torque on the binary from the gas parcels that are distributed in the disk. The actual efficiency of mass accretion onto stellar-mass BBH from disks formed in post-CE systems is influenced by multiple uncertain factors, e.g., magnetic fields in gas disks of this type, which affect the rate of viscous transport.
We consider three accretion paradigms in this work. The first is unlimited accretion, representing a binary in an infinite gas reservoir that continually accretes mass. The second is no accretion ("none" in Fig. 2), representing a binary in an infinite gas disk that continually interacts with the disk but does not accrete any matter at all. Negligible accretion may be the norm for post-CE BBHs with disks: “typical” binary-disk parameters for post-CE stellar-mass BBH are found to converge towards the zero-accretion case due to the torques that the binary exerts on the disk (Martin et al., 2018). The third case is limited accretion, in which the disk mass is 10% of $M_{i}$, and the binary accretes until the disk runs out.
@@ 3 Results
Stellar-mass BBHs with CBDs are candidates for multi-band observations and archival searches in LISA data. Fig. 2 shows the expected eccentricity at $0.01$ Hz, $e_{0.01}$. We find $e_{0.01}\sim 10^{-2}$, within the range of detectability for such searches, and of the same magnitude as expected eccentricities from dynamically-formed mergers (Wang et al., 2024). The $e_{0.01}$ values remain set primarily by $q_{i}$, although for the unlimited-accretion case we expect $q_{f}\approx 1$. For the negligible-accretion case, in which $q_{f}=q_{i}$, this correlation a potential smoking-gun of BBHs with CBDs. We note that all $e_{0.01}$ values are given in the source frame, without considering the effects of the redshifting of the reference frequency. Using Peters (1964), we estimate that an eccentricity of $e=3\times 10^{-2}$ at $0.01$ Hz is redshifted to $e=2\times 10^{-2}$ if $z=0.2$ ($10^{-3}$ if $z=1$); see also Sec. 4.1.1 of Romero-Shaw et al. (2022). Redshifting therefore likely accounts for the much lower predicted $e_{0.01}\sim 2\times 10^{-3}$ for the $q_{i}=1$ BBH with $M_{i}=147$ M highlighted in Zrake et al. (2021), although higher-mass BBHs also circulate at lower frequencies than lower-mass BBHs.
@@
By the time the $f_{\mathrm{GW}}$ reaches $11$ Hz, the optimistic minimum frequency of the future ground-based detector Einstein Telescope (ET; Maggiore et al., 2020), BBHs with CBDs have eccentricities $e_{1}\approx 10^{-4}$ if $f_{e}=0.1$, equivalent to $e_{10}\approx 10^{-5}$, below the expected sensitivity for ET (Lower et al., 2018; Saini, 2024).
@@ 4 Discussion
The LISA-band eccentricity evolution of a stellar-mass BBH is markedly different when it has a CBD, as the CBD drives the binary to a characteristic equilibrium eccentricity determined by its initial mass ratio. We find that both $e_{\rm eq}$ and $e_{0.01}$ directly correspond to $q_{i}$, insensitively to assumptions about accretion efficiency. As long as the disk endures, the eccentricity at $f_{\mathrm{GW}}=0.01$ Hz is $e_{0.01}\sim 10^{-2}$, detectable for LISA and in the range usually expected for diskless dynamically-formed in-cluster mergers (Klein et al., 2022; Garg et al., 2023; Wang et al., 2024; DePorzio et al., 2024). Since $e_{0.01}$ is a strong function of $q_{i}$, a population of BBHs observed with an eccentricity-mass ratio correlation would indicate contributions from non-accreting BBHs with CBDs, while a population of $q=1$ stellar-mass BBHs with eccentricities $e_{0.01}\sim 2\times(10^{-3}$$10^{-2})$ would more likely come from accreting BBHs with CBDs. The values of $e_{0.01}$ that we predict for stellar-mass BBHs with CBDs are similar to those predicted for more massive BBHs with CBDs when they enter the LISA band (Armitage & Natarajan, 2005; Roedig et al., 2011; Zrake et al., 2021; Siwek et al., 2024); across all mass scales, BBHs with CBDs may be characterised by distinct mass ratios and eccentricities in LISA, once redshifting is corrected for.
@@
The eccentricity evolution of stellar-mass BBHs with CBDs is similar to that of the supermassive BBHs evolved in Siwek et al. (2024), but proceeds at higher frequencies and slightly higher eccentricities due to the lower mass. Our results using limited accretion agree with those of Valli et al. (2024): disks with $M_{d}=0.1M_{i}$ do not drastically influence the separation of the BBH, but do drive the binary eccentricity to its equilibrium value. The $e_{\mathrm{eq}}$ values we find for stellar-mass BBHs are in good agreement with the equilibrium eccentricity values found in Valli et al. (2024), with the maximum equilibrium eccentricity of $e_{\mathrm{eq}}\approx 0.5$ occurring for $q_{i}=0.7$. However, in contrast to previous work including Siwek et al. (2023b, a); Valli et al. (2024), none of our binaries undergoes orbital expansion, since $\dot{a}_{\mathrm{GW}}\gg\dot{a}_{\mathrm{gas}}$ when $a_{i}=0.2$ au.